A parachutist after bailing out falls 50m without friction when it opens it deaccelerates with 2m/s2. he raeches the ground with 3m/s at wat height did he ball out?

Grade:

1 Answers

Farhad Nawaz
11 years ago

(downward as +ve) initially u=0, after 50m, let velocity be V. for 50m the parachute undegoes freefall. 50=V^2/2g. V=10root(10) m/s. for the rest of the journey, a=-2m/s^2, at ground level, u=3m/s. for this journey alone, let the drop in height be H. u^2-V^2=2aH. H=247.75m=answer.